OpenStudy (anonymous):

find the equation of the line tangent to 9x^2+16y^2=52 through (2,-1)

OpenStudy (anonymous):

take the derivative wrt x and get \[18x+32yy'=0\] solve for \(y'\) and get \[y'=-\frac{18x}{32y}=-\frac{9x}{16y}\] replace x by 2 and y by -1 to get the slope, then use the point - slope formula

OpenStudy (anonymous):

ohhh gotcha. thanks

OpenStudy (anonymous):

yw